records anuary February 91,800 $144,e0e 70,se0 .The 31, 20x0, balance sheet revealed the following selected figures: cash. $24,000; accounts receivable, $105,000: and finished goods, $24,600 Required: the number of units that Badlands sold in December 20x0 20x1 to be 31, 20x0.
Answer #1

1)The collections on December sales will occur as follow: 30% in December; 70% in January.

$105,000 was collected in January from December sales; this represents represents 70% of December sales.

So December Sales= $105,000/70%= 150,000

In units= 150,000/25= 6,000 Units

2)Expected March sales Revenue= 7,000 units x $25/unit = $175,000

3)

Description Amount
January sales ($91,000+ $144,000) $ 235,000.00
February sales ($70,500 ÷ 30%) $ 235,000.00
March sales (from #2 – 7,000 units x $25 unit) $ 175,000.00
Total Sales revenue $ 645,000.00

4)70% of March sales will be collected in April. Therefore, the March 31 Accounts receivable balance is $175,000 x 70% =$122,500

5) Finished goods inventory = 30% of the following month’s sales

January sales = $235,000 ÷ $25/unit = 9,400 units

December 31 finished goods inventory = 30% x 9,400 = 2,820 units

6)

January 1 finished goods inventory   $      2,820.00
January 31 finished goods inventory 30% x ($235,000 ÷ $25) =
$      2,820.00
January Sales $      9,400.00
Finished goods to be manufactured in January $      9,400.00

7)

Cash balance, January 1 $                             24,000.00
Add :January Receipts $                           196,000.00
Deduct:  January payments $                        (205,000.00)
Cash balance, January 31 $                             15,000.00
Desired minimum balance   $                           (21,000.00)
Required financing   $                             (6,000.00)